Bay State Community College enrolls students in two departments, Liberal Arts and Sciences. The college also has two service departments, the Library and the Computing Services Department. The usage of these two service departments’ output for the year is as follows:

Provider of Service
  
  User of Service Library Computing Services
  Library 10 %
  Computing Services
  Liberal Arts 80 % 20 %
  Sciences 20 % 70 %
The budgeted costs in the two service departments for the year are as follows:
  
  Library $ 730,000
  Computing Services 360,000
Required:
1.

Use the direct method to allocate the budgeted costs of the Library and Computing Services Department to the college’s Liberal Arts and Sciences departments. (Do not round intermediate calculations.)

Library

Computing

Total
Answer #1
Library Computing Services Liberal Arts Sciences
Departmental Cost of Library and Computing Services $ 730,000 $ 360,000                                     -                                   -  
Allocation :
Library Department Cost $ ( 730,000 )                                        -   $ 584,000 $ 146,000
Computing Services Cost $ ( 360,000 ) $ 80,000 $ 280,000
Total $ 664,000 $ 426,000
Library Department Cost = $ 730,000
Allocation :
Liberal Arts : Science = 80 % : 20%
Liberal Arts = $ 730,000 * 80 % = $ 584,000
Science = $ 730,000 * 20% = $ 146,000
Computing Services Department Cost = $ 360,000
Allocation :
Liberal Arts : Science = 20 % : 70%
Liberal Arts = $ 360,000 * ( 2 / 9 ) = $ 80,000
Science = $ 360,000 * ( 7 / 9 ) = $ 280,000